OpenStudy (anonymous):

Differentiate ln(2x+4)(3x-1)

Parth (parthkohli):

\[\ln[(2x + 4)(3x - 1)]\]?

Parth (parthkohli):

If that is the case, then use the Chain Rule.

OpenStudy (anonymous):

does it help to know that this is \[\ln(2x+4)+\ln(3x-1)\]? then it should much easier